OpenStudy (anonymous):

What is the simplified form of 2x + 4 over x - 2 minus x-3 over x-2

sam (.sam.):

\[\frac{2x + 4}{x-2}-\frac{x-3}{x-2}\] since the denominator is same, they can be combined \[\frac{(2x + 4)-(x-3)}{x-2}\]

sam (.sam.):

then just do the numerator math \[\frac{(2x + 4)-(x-3)}{x-2}\] \[\frac{2x + 4-x+3}{x-2}\] \[\frac{x+7}{x-2}\]